solve by substitution method
2x-9y= -8
9x+115=y what is the solution set. I came up with (17, -9) am I close

Answer by checkley77(12844)

2x-9y=-8
9x-Y=-115 MULTUPLY THIS EQUATION BY -9 & ADD.
-81X+9Y=1,035
2X-9Y=-8
--------------
-79X=1,027
X=1,027/-79
X=-13 ANS.
2*-13-9Y=-8
-26-9Y=-8
-9Y=-8+26
-9Y=18
Y=18/-9
Y=-2 ANS.
PROOF:
9*-13-(-2)=-115
-117+2=115
-115=-115